What is the sum of money to be accumulated by each $5000 investment in 31 years and compound at 7% per year. And $20000 in 7 years after compounding 6% annually.
Answer by Boreal(15235)

P=Po(1+r)^n, when this is annually. Po is start, r is rate,n is year.
=5000(1.07)^31=$40,725.56
======================
=20000(1.06)^7=
$30,072.61
round everything at the end.